Step 1: You tossed a coin 50 times and got 21 heads. The proportion of heads...

Step 1:
You tossed a coin 50 times and got 21 heads. The proportion of heads is pˆ= 21/50 = 0.42. The proportion is less than 0.5. You want to find out whether this is evidence that your coin is not balanced.
Step 2:
What conclusion can you make about this coin?

o Because the chance of observing 21 heads in 50 tosses is large, we do not reject H 0 and conclude the coin is balanced.

o Because the chance of observing 21 heads in 50 tosses is small, we reject H 0 and conclude the coin is balanced.

o Because the chance of observing 21 heads in 50 tosses is large, we reject H 0 and conclude the coin is not balanced.
Answer #1

​​​​​​To test

Ho: Po=0.5,. V/S. H1:Po0.5

p^=21/50= 0.42 ,n=50

Test statistic

Z= P^-Po/

Z=0.42-0.50/

Z=-1.1313

Critical value for z test at alpha level 0.05 is -1.64

Where,. Z cal > Z critical

We accept Ho

#Conclusion: the chance of getting 21 heads in 50 tosses is large , we do not reject Ho and conclude the coin is balanced